# TikTokActions: A TikTok-Derived Video Dataset for Human Action Recognition

### v2

@@ Abstract
The increasing variety and quantity of tagged multimedia content on platforms such as TikTok provides an opportunity to advance computer vision modeling. We have curated a distinctive dataset of 283,582 unique video clips categorized under 386 hashtags relating to modern human actions. We release this dataset as a valuable resource for building domain-specific foundation models for human movement modeling tasks such as action recognition. To validate this dataset, which we name TikTokActions, we perform two sets of experiments. First, we pretrain the state-of-the-art VideoMAEv2 with a ViT-base backbone on TikTokActions subset, and then fine-tune and evaluate on popular datasets such as UCF101 and the HMDB51. We find that the performance of the model pre-trained using our TikTok dataset is comparable to models trained on larger action recognition datasets(95.3% on UCF101 and 53.24% on HMDB51). Furthermore, our investigation into the relationship between pre-training dataset size and fine-tuning performance reveals that beyond a certain threshold, the incremental benefit of larger training sets diminishes. This work introduces a useful TikTok video dataset that is available for public use and provides insights into the marginal benefit of increasing pre-training dataset sizes for video-based foundation models.
@@ 1 Introduction
Action recognition is a fundamental computer vision task that involves identifying and classifying human actions in video sequences, with significant applications in areas such as surveillance, interactive gaming, and healthcare. In the era of social media platforms, TikTok has emerged as a significant repository of human action videos. We have compiled a dataset of 283,582 TikTok video clips that encompass a range of human activities, including dance, sports, fitness, hand gestures, and daily routines. This collection aims to represent various human actions in a manner that is relevant to real-world scenarios and cultural contexts. Importantly, this dataset is uniquely curated to include well-defined but non-standard human actions, filling gaps in existing action recognition datasets and diversifying the range of human activities recognized by current models. We developed a model for human action recognition trained using the curated TikTok video clips. Pre-training on a dataset with varied and authentic content is intended to more accurately reflect the spectrum of naturalistic human actions, aiming to address the limitations of conventional datasets by incorporating elements of contemporary behavior. In our initial evaluations, presented later in this paper, we find that utilizing our TikTok dataset may enhance the development of targeted human action recognition models in the field of computer vision.
Moreover, the nature of our data source, TikTok, presents another advantage. Kinetics[4] derives its content from YouTube, featuring a combination of professional and amateur videos, whereas TikTok’s database consists primarily of user-generated material. This means that our dataset is inherently rich with real-world scenarios: from spontaneous camera movements and varied illumination conditions to diverse backgrounds and authentic human interactions. This rawness provides a genuine challenge and opportunity for training robust models capable of understanding human actions in a plethora of real-world conditions.
We conducted a series of experiments to assess the quality of our dataset and its potential to serve as a basis for pre-training foundation models for action-related tasks. We used a well-established state-of-the-art model VideoMAE V2[9], and pre-trained it on our TikTokActions dataset. We then fine-tuned the model on two well-known video action recognition datasets: UCF101[7] and HMDB51[5]. Our experiments provided valuable insights: the performance achieved was comparable to models that were both pre-trained and fine-tuned solely on UCF101 and HMDB51. This underscores the utility of the TikTokActions dataset to serve as a useful pre-training dataset for action-related tasks.
We next delved into understanding the relationship between the quantity of pre-training videos and the performance on the downstream action recognition task. By sampling subsets of our TikTok dataset, starting with 1,000 videos and increasing in increments of 1,000 up to 6,000 videos, pre-training VideoMAE V2[9] on the sampled subset, and subsequently fine-tuning each on the benchmark datasets, we observed that as long as the number of pre-training videos remains above a certain threshold, the performance remains comparable to massive pre-training datasets. This provides a valuable lesson when pre-training foundation models: the benefits of increasing the pre-training dataset size diminish as the dataset increases.
@@ 3 Dataset Construction
3.1 Dataset Overview
We introduce the TikTokActions dataset, designed for human action recognition. Derived from TikTok videos, this dataset comprises over 280,000 video clips from 386 hashtags that capture a wide spectrum of human actions in various scenarios. The dataset spans diverse categories, from dance movements and fitness exercises to hand gestures and daily activities. The aim of TikTokActions is to provide a robust and comprehensive foundation for training and evaluating human action recognition models in realistic settings.
3.2 How the Dataset was Built
To identify a list of hashtags to mine on TikTok, we initially consulted several commonly used datasets related to human actions, including UCF101[7], HMDB51[5], and Google DeepMind Kinetics[4]. After conducting a thorough comparison of these datasets, we identified four categories of videos that can best represent various human actions: dance, sports, fitness, and kinetics. Following these categorizations, we proceeded to identify several hashtags under each category. During the process of structuring our hashtag list, we drew insights from the hashtags commonly employed by TikTok users on the platform. This comparative analysis yielded six fundamental principles that guided our approach to selecting hashtags to mine for data: generalization, specification, noun substitution, noun-verb order, hashtags consolidation, and social/cultural consideration. The last principle enabled our team to include the fifth category unique to social media, encompassing a wide variety of challenge types of actions originating from distinct social and cultural contexts. Additionally, beyond these six principles, we placed significant emphasis on the number of views under each hashtag, recognizing that popularity and engagement levels are crucial indicators of the relevance and impact of the content for each hashtag. We set a 5 million view minimum threshold for hashtag inclusion, reflecting a focus on content demonstrating substantial public engagement and visibility, indicative of broad influence.
@@ 3.2.3 Noun Substitution
Table 1: Table of Hashtags and Their Corresponding Number of Views
Figure 1: Flowchart depicting the hashtag selection and modification process.
@@ 3.2.6 Social/cultural Considerations
3.3 Data Collection and Processing
We constructed the TikTokActions dataset using the official TikTok research API. For each hashtag, approximately 900 videos and corresponding metadata were mined. Each video underwent processing with tools like PySceneDetect and YOLO v8[6] producing a JSON file for every video, thereby accumulating metadata for 340,000 videos. The metadata encompassed various attributes, including the title, duration, and scene distribution based on PySceneDetect. We describe details about the preprocessing procedure below.
3.3.1 PySceneDetect and YOLOv8 Preprocessing
Our pre-processing approach began with the use of PySceneDetect to analyze the scene composition of each video. PySceneDetect was employed to identify distinct scenes within the videos, enabling us to focus on the longest scene clips. The rationale behind selecting the longest scenes was our hypothesis that these segments have the highest probability of containing human presence.
Once the longest scenes were extracted, we then utilized the YOLOv8[6] model to detect human figures in these video segments. The primary goal of this method was to accurately identify videos containing human subjects. The procedure comprised several distinct phases, each fulfilling a specific aspect of the preprocessing objective.
In the frame extraction phase, the total frame count of each video clip was determined using the ffprobe tool. Given the considerable frame count in many videos, it was computationally impractical to process each frame. Consequently, we resorted to subsampling. Typically, a fixed number of frames, defaulting to 10, were extracted at regular intervals. For a video containing 101 frames, every 10th frame was extracted starting from the 10th frame.
The next step was human detection. Here, the extracted frames underwent inference using the YOLOv8[6] model, initialized based on Ultralytics’ YOLO framework. The primary task was to review the detections and identify frames with human presence. Frames with at least one ”person” label contributed to the cumulative count of frames with human subjects for that particular video.
Lastly, the post-processing steps included thresholding. For a video to qualify as containing humans, it was required to surpass a pre-defined frame threshold with human detections, set to 4 by default. Videos below this threshold were earmarked for subsequent human scrutiny. Following the analysis of all videos, a results compilation phase calculated aggregated statistics, providing a clear picture of the number of videos with humans. Additionally, any extracted frames were systematically removed post-inference to maintain a clean working environment. In Figure  2, we show the process of the generation of clips for each video.
Figure 2: Flowchart depicting the clip extraction process from raw videos mined on TikTok.
@@ 4 Performance on Benchmarks
In the domain of human action recognition, benchmarking performance is crucial to understanding the efficacy of a model. In this section, we evaluate the performance of our model, which was initially pre-trained on VideoMAE V2[9] with TikTokActions subset and subsequently fine-tuned on two datasets: UCF-101[7] and HMDB51[5].
@@ 4.1 Pre-training on VideoMAE V2
VideoMAE V2[9] represents a state-of-the-art masked autoencoding framework with a billion-level ViT backbone. This framework has been instrumental in enhancing performance across a diverse range of video-related tasks, including video action recognition, action detection, and temporal action detection. The generalization ability of VideoMAE V2 pre-trained ViTs[1] as video foundation models has been verified across diverse downstream tasks, showcasing its robustness and adaptability.
Given the strength of VideoMAE V2[9], we utilized it for our pre-training phase. Table 2 displays the hyperparameters used for this phase. The weights of VideoMAE V2 were randomly initialized.
Table 2: Hyperparameters for Pre-training on VideoMAE V2
Parameter Value
Mask Ratio 0.9
Decoder Mask Ratio 0.5
Model Vit_base
Decoder Depth 4
Batch Size 32
Number of Frames 16
Sampling Rate 4
Number of Samples 2
Number of Workers 24
Optimizer adamw
Learning Rate 1e-3
Optimizer Betas 0.9, 0.95
Warmup Epochs 20
Total Epochs 200
We conducted our training on a single NVIDIA A100 40GB GPU. This constraint necessitated a reduction in the scale of our dataset; consequently, we focused on a randomly selected subset comprising 10,000 videos, selected from the total collection of 280,000. This approach allowed us to manage the computational demands while still providing a representative sample for training and evaluating the model’s performance. This constraint also enabled us to study the utility of self-supervised pre-training on smaller datasets and compare it to prior work which are pre-trained on larger datasets.
@@ 4.2 Fine-Tuning on Established Datasets
Parameter Value
Batch Size 8
Input Size 224
Short Side Size 224
Number of Frames 16
Sampling Rate 4
Number of Workers 24
Optimizer adamw
Learning Rate 1e-3
Drop Path 0.35
Layer Decay 0.9
Optimizer Betas 0.9, 0.999
Weight Decay 0.05
Total Epochs 100
Test Number of Segments 5
Test Number of Crops 3
Table 3: Hyperparameters for Fine-tuning
@@
HMDB-51[5] 53.24 89.73
UCF-101[7] 95.3 99.54
Table 4: Fine-tuning Results With 10,000 Tiktok Actions videos Pre-training
After the initial pre-training on VideoMAE V2[9], the model was fine-tuned on established datasets. The hyperparameters for the fine-tuning phase are provided in Table 3. The results of the fine-tuning on various datasets, UCF-101[7] and HMDB51[5] are consolidated in Table 4. It showcases the effectiveness of the TikTok Actions dataset in model pre-training. Notably, the model achieved a high Top-1 accuracy of 95.3% on UCF-101 and 53.24% on HMDB-51, demonstrating its strong generalization capabilities across diverse action recognition tasks.
@@ 5 Extrapolating the Required Pre-Training Dataset Size
t
9157 12388 18323 31892 82574
Burr XII
9090 12209 17857 30511 76240
InvGamma
9072 12166 17750 30204 74859
BetaPrime
9072 12165 17748 30200 74843
F
9072 12166 17749 30203 74855
GenPareto
9072 12164 17744 30184 74759
Lomax
9072 12164 17744 30184 74759
LogLaplace
9072 12164 17744 30184 74760
Pareto
9072 12164 17743 30183 74755
InvWeibull
9043 12101 17596 29772 72915
@@ 6 Discussion
The development of the TikTok-sourced dataset aimed to contribute to the action recognition field by including a variety of well-defined but non-standard human actions. These actions were selectively curated to complement the actions typically found in existing datasets, rather than duplicating them. This approach in dataset composition was chosen to diversify the range of human actions available for training and evaluating action recognition models.
By integrating these unique and diverse actions from the TikTok platform, our dataset broadens the scope of actions recognizable by machine learning models. This expansion is not merely a quantitative increase in data but a qualitative enrichment that challenges and enhances the adaptability and comprehensiveness of action recognition algorithms. This contribution is particularly significant in light of the evolving nature of human actions in digital media, where new and unconventional actions frequently emerge.
We also made auxiliary findings with notable implications: we observed a diminished necessity for large pre-training dataset sizes. Even with a relatively smaller subset of the TikTokActions dataset, the model performance remains competitive. This challenges the prevailing notion in action recognition research that larger datasets are invariably superior for pre-training. The implications of this finding indicate a potential shift in the paradigm of dataset compilation and model training. It points towards the efficiency of a smaller, well-curated dataset that can significantly reduce computational resources while maintaining high accuracy, which may be preferred over a larger but less valuable dataset.
One of the primary limitations of this study is the incomplete pre-training on the full TikTokActions dataset. Due to computational constraints, our experiments were conducted on a subset of 10,000 videos, which might not fully represent the diversity and complexity of the full dataset. Additionally, while the dataset performed well on benchmarks like UCF101[9] and HMDB51[5], its efficacy across other popular action recognition datasets remains untested.
Another significant limitation is the absence of ground truth labels for each clip in the TikTokActions dataset. Although the videos were curated scientifically, the lack of verified labels poses a challenge for more rigorous supervised learning approaches, potentially impacting the precision of the training and evaluation processes.
Future research endeavors can utilize the entirety of or subsets of the TikTokActions dataset for pre-training, aiming to enhance model performance and generalization across action recognition tasks. Another opportunity is the integration of TikTokActions with other datasets such as Kinetics[4], which will provide insights into the unique features and potential synergies of these datasets. This comparison is crucial for validating TikTokActions’ effectiveness and identifying areas for expansion or enhancement.
@@
An emerging area of exploration is the application of weekly self-supervised learning techniques. These methods, which involve regular, unsupervised updates to the model based on new data influxes, could significantly enhance the adaptability and accuracy of the models trained on TikTokActions.
Furthermore, employing crowd-sourcing to acquire ground truth labels for the TikTokActions dataset presents a promising opportunity. This initiative would enable researchers to conduct more supervised learning studies, potentially leading to more accurate and reliable models. By combining crowd-sourced annotations with sophisticated learning algorithms, the field can expect to see substantial advancements in action recognition capabilities and applications.

### v3

# Advancing Human Action Recognition with Foundation Models trained on Unlabeled Public Videos

@@ Abstract
The increasing variety and quantity of tagged multimedia content on a variety of online platforms offers a unique opportunity to advance the field of human action recognition. In this study, we utilize 283,582 unique, unlabeled TikTok video clips, categorized into 386 hashtags, to train a domain-specific foundation model for action recognition. We employ VideoMAEv2, an advanced model integrating Masked Autoencoders (MAE) with Vision Transformers (ViT), pre-trained on this diverse collection of unstructured videos. Our model, fine-tuned on established action recognition benchmarks such as UCF101 and HMDB51, achieves results comparable with the state-of-the-art: 97.5% on UCF101, 83.9% on HMDB51, 81.1% on Kinetics-400, and 69.2% on Something-Something V2 using the ViT-base backbone. These results highlight the potential of using TikTok videos as a valuable source of diverse and dynamic content for training foundation models. Our investigation confirms that while initial increases in pre-training data volume significantly enhance model performance, the gains diminish as the dataset size continues to expand. Our findings identify two critical lessons in self-supervised learning for computer vision: (1) additional pre-training data yields diminishing benefits and (2) these diminishing benefits can be observed at relatively small dataset sizes on the order of thousands of images.
@@ 1 Introduction
Action recognition is a fundamental computer vision task that involves identifying and classifying human actions in video sequences, with significant applications in areas such as security, interactive gaming, and healthcare. In the era of social media platforms, TikTok has emerged as a significant source of human action videos. We utilized a collection of 283,582 TikTok video clips, encompassing a wide range of human activities, to train a foundation model tailored to human action recognition. This internally curated collection aims to represent various human actions relevant to real-world scenarios and cultural contexts. Importantly, it includes well-defined but non-standard human actions, addressing gaps in existing action recognition datasets and diversifying the range of human activities recognized by current models.
Unlike Kinetics [4], which also includes user-generated content from YouTube, our collection exclusively features TikTok videos, known for their dynamic and spontaneous nature. This distinction offers unique training advantages, as such videos often exhibit quick transitions, novel perspectives, and creative interactions—elements less prevalent in YouTube’s often longer-form content. These characteristics provide robust challenges for training models, enhancing their ability to interpret a wide range of human actions in diverse and unpredictable environments.
We conducted a series of experiments to assess the potential of videos collected in this manner for pre-training computer vision foundation models for action-related tasks. We employed the well-established, state-of-the-art model VideoMAE V2, using it with ViT-base backbones[9]. The ViT-base was used to provide a baseline comparison with other models using similarly sized backbones. Subsequent fine-tuning on well-known video action recognition datasets like UCF101[7], HMDB51[5], Kinetics[4], and Something-Something V2[3] demonstrated valuable insights: the performance achieved with both backbones was comparable to, or in some cases even better than, models trained exclusively on these conventional datasets or even those pre-trained on much larger datasets, where performance increases are often marginal.
We also explored the relationship between the quantity of pre-training videos and the performance on downstream tasks. By sampling subsets of our collection, starting with 1,000 videos and increasing in increments of 1,000 up to 6,000 videos, and pre-training VideoMAE V2[9] on these subsets, we observed that performance remains robust as long as the number of pre-training videos stays above a certain threshold, highlighting that massive datasets may not always be necessary for effective pre-training. This exploration not only emphasizes the efficiency of using online unlabeled videos as a pre-training source but also challenges the conventional reliance on extensive video datasets for achieving high model performance. Before deployment of this approach in the real world, we encourage researchers to think through the ethics of data scraping from publicly available sources to train deep learning models.
@@ 3 Dataset Construction
3.1 Video Collection Overview
To support our experimentation exploring the utility of unlabeled public data for pre-training foundation models for human action recognition, we curated an extensive collection of TikTok videos. This collection, consisting of over 280,000 video clips from 386 hashtags, captures a wide spectrum of human actions across various scenarios. The videos span diverse categories, from dance movements and fitness exercises to hand gestures and daily activities. Unlike traditional datasets like Kinetics, our collection uniquely incorporates elements of pop culture, memes, and rapid trend evolution, providing dynamic content that significantly boosts the model’s ability to understand and generalize from real-world, culturally relevant scenarios.
3.2 Video Selection Methodology
To curate a diverse and relevant set of videos from TikTok, we initially examined several well-known action recognition datasets, including UCF101[7], HMDB51[5], and Google DeepMind Kinetics[4]. These datasets helped us to understand the broad categories of human actions typically represented in research. We identified four primary categories: dance, sports, fitness, and kinetics.
Building upon these categories, we explored TikTok to identify relevant hashtags that capture a wide array of human actions within these domains. Our selection was strategically guided by six principles designed to ensure a comprehensive and culturally relevant collection: generalization, specification, noun substitution, noun-verb order, hashtag consolidation, and social/cultural relevance. These principles helped ensure that the videos not only cover a broad spectrum of actions but also include dynamic and culturally significant content such as memes, pop culture trends, and unique challenges—elements that are distinctly prevalent on TikTok.
The social/cultural relevance was particularly important, allowing us to incorporate an additional category that includes a wide variety of actions originating from distinct social and cultural contexts, such as viral dance challenges and fitness trends. Popularity and engagement levels also played a crucial role in our selection process; we set a minimum threshold of 5 million views per hashtag to ensure that the content is not only relevant but also resonates widely with users, reflecting substantial public engagement and visibility. This approach allows us to harness rich, culturally nuanced content that can aid in training models to better understand and interpret human actions in real-world contexts.
@@ 3.2.3 Noun Substitution
Table 1: Example of the expansion in data availability when removing nouns from the hashtag construction process.
Figure 1: The hashtag selection and modification process.
@@ 3.2.6 Social/cultural Considerations
3.3 Video Collection and Processing
We curated a comprehensive video collection using the official TikTok research API, focusing on approximately 900 videos per hashtag from a set of hashtags identified through our selection process. Each video and its metadata were processed to facilitate model training, using tools such as PySceneDetect and YOLO v8[6]. This processing resulted in a JSON file for each video, which included metadata capturing attributes like title, duration, and scene distribution, as analyzed by PySceneDetect.
3.3.1 PySceneDetect and YOLOv8 Processing
Our processing approach began with PySceneDetect, used to analyze the scene composition of each video. This tool helped identify distinct scenes within the videos, enabling us to focus on the longest scene clips, hypothesized to have the highest likelihood of containing significant human action.
Once the longest scenes were identified, the YOLOv8[6] model was employed to detect human figures in these segments. The primary goal of this processing was to ensure accurate identification of videos containing human subjects, vital for the training of our models.
During the frame extraction phase, we used the ffprobe tool to determine the total frame count of each video clip. Due to the large number of frames in many videos, we applied subsampling for efficiency. Typically, a fixed number of frames (usually 10) were extracted at regular intervals throughout the video.
The extracted frames were then processed using the YOLOv8[6] model to detect human presence. Frames that included at least one ”person” label were counted towards the total number of frames containing humans for that particular video.
Post-processing included applying a threshold for human presence; videos needed to exceed a pre-defined frame threshold with human detections to qualify. Videos that did not meet this threshold were subject to further review. This systematic approach helped compile aggregated statistics on the human presence in our video collection, ensuring the quality of content used for model training. The entire process, from video selection to frame extraction and human detection, was documented to maintain transparency and reproducibility. Figure  2 illustrates the process of clip extraction from the raw videos sourced from TikTok.
Figure 2: The clip extraction process from raw videos mined on TikTok.
@@ 4 Performance on Benchmarks
In the domain of human action recognition, benchmarking performance is essential to understand the efficacy of a model. This section evaluates the performance of our model, initially pre-trained on the entire collection of 283,582 unlabeled TikTok video clips using VideoMAE V2[9], and subsequently fine-tuned on Kinetics-400 (K400). Following this, the model underwent further fine-tuning on UCF-101[7] and HMDB51[5], adopting a transfer learning approach akin to the experimental setups used in VideoMAE V2 studies. Additionally, we extended our evaluation to include the Something-Something V2[3] dataset, assessing the model’s adaptability and performance across varied action recognition contexts.
@@ 4.1 Pre-training on VideoMAE V2
VideoMAE V2[9] is a state-of-the-art masked autoencoding framework, notable for its flexibility in backbone selection, ranging from base to billion-level configurations. We opted for the ViT-base backbone, which, while less computationally intense than its billion-level counterparts, significantly enhances performance across various video-related tasks, including video action recognition. The generalization capability of VideoMAE V2 pre-trained ViTs[1] as video foundation models has been demonstrated across multiple downstream tasks, highlighting its robustness and adaptability.
Our training was conducted on 8 NVIDIA A100 40GB GPUs, utilizing the full set of 283,582 unlabeled TikTok videos. The pre-training on the ViT-base backbone lasted for 20 days, providing a comprehensive base for developing a robust model capable of understanding a wide array of human actions.
@@ 4.2 Fine-Tuning on Established Datasets
After initial pre-training, the model underwent two distinct paths of fine-tuning. The first path involved direct fine-tuning on the Kinetics-400 and Something-Something V2 datasets to assess immediate applicability to broad action recognition tasks. The results of this direct fine-tuning are detailed in Table 2. The second path involved using the model fine-tuned on Kinetics-400 as an intermediary for transfer learning on UCF-101 and HMDB51, leveraging robust pre-training on TikTok videos. The outcomes of this transfer learning process are presented in Table 3.
@@
Kinetics-400[4] 81.1 94.8
Something-Something V2[3] 69.2 91.6
Table 2: Direct Fine-Tuning Results With TikTok Videos Pre-training
Dataset Top-1 (%) Top-5 (%)
HMDB-51[5] 83.9 96.7
UCF-101[7] 97.5 99.6
Table 3: Transfer Learning Results From K400 Fine-Tuning Through TikTok Pre-training
The fine-tuning results highlight the significant potential of using large, unlabeled video collections from platforms like TikTok for advancing human action recognition tasks. This approach capitalizes on the rich, varied nature of content available on social media, which, when used for pre-training, enhances the model’s ability to generalize across different datasets and recognition scenarios. The robust performance across both direct and transfer learning paths underscores the value of TikTok videos in training models to tackle complex action recognition challenges.
@@ 5 Extrapolating the Required Pre-Training Dataset Size
t 9157 12388 18323 31892 82574
Burr XII 9090 12209 17857 30511 76240
InvGamma 9072 12166 17750 30204 74859
BetaPrime 9072 12165 17748 30200 74843
F 9072 12166 17749 30203 74855
GenPareto 9072 12164 17744 30184 74759
Lomax 9072 12164 17744 30184 74759
LogLaplace 9072 12164 17744 30184 74760
Pareto 9072 12164 17743 30183 74755
InvWeibull 9043 12101 17596 29772 72915
@@ 6 Discussion
The curation of TikTok videos for enhancing human action recognition models provides a unique contribution to the field, especially by incorporating a variety of distinct yet non-traditional human actions. These actions, sourced from the TikTok platform, were carefully selected to complement the more conventional actions typically found in established datasets like UCF101[7] and HMDB51[5]. This approach not only diversifies the types of human actions available for model training but also qualitatively enriches the dataset, thereby challenging and extending the adaptability of action recognition algorithms.
Integrating these unique and diverse actions from TikTok, we have expanded the range of actions that machine learning models can recognize, addressing the dynamic nature of human actions in digital media where new and unconventional actions frequently emerge. This strategy highlights the value of leveraging culturally rich and dynamic user-generated content as a training resource in an era dominated by digital media.
Furthermore, our findings challenge the conventional wisdom that larger datasets always yield better pre-training results. The competitive performance of our models, initially pre-trained on a curated subset of TikTok videos and subsequently fine-tuned on Kinetics-400[4] (K400) and Something-Something V2(Sth-Sth)[3], suggests that a well-curated, smaller dataset can sometimes outperform a larger, more generic one. This observation encourages a shift towards more efficient dataset compilation and model training paradigms, emphasizing data quality over quantity.
Looking ahead, exploring the potential of using the full breadth of the TikTok video collection for pre-training could further enhance model performance and generalization across various tasks. Additionally, integrating these videos with other datasets, such as Kinetics-400 and Something-Something, could provide deeper insights into the unique characteristics and potential synergies between different types of video content.
@@
An emerging area of interest is the application of weekly self-supervised learning techniques, which involve regular, unsupervised updates to the model based on new data inflows. Such techniques could significantly boost the adaptability and accuracy of models trained with dynamic, culturally relevant content from platforms like TikTok.
7 Ethics
This research adheres to strict ethical guidelines regarding the use of online video data, specifically videos collected from TikTok. All videos were legally accessed via the TikTok Research API, and only those videos where the user granted download permission were included. Our data collection process is in full compliance with the TikTok Research API Terms of Service. As an additional layer of security, we delete the data after training.
We place a high priority on privacy and copyright issues. To protect individual privacy, we have implemented face obfuscation techniques in the video clips included in our paper. Moreover, in line with the terms of service, we do not distribute any TikTok user content directly. Instead, we focus on training a foundation model, which we only use to identify action patterns, not personal biometric features.
Ethical Considerations in Online Video Research: The use of online video data poses unique challenges and ethical considerations, particularly regarding privacy, informed consent, and the potential risks associated with the use of such data. Our study emphasizes the importance of adhering to ethical principles that safeguard participant interests. We have made significant efforts to minimize potential harm by anonymizing the data and not sharing the scraped dataset. Although techniques like pixelating faces can be circumvented and are not foolproof, they represent our commitment to reducing identifiable information.
Furthermore, we advocate for the inclusion of clearer guidelines in the terms of service on video platforms, which would allow users to explicitly exclude their videos from research use. This approach does not replace the need for informed consent but is a step towards enhancing ethical practices in online video research. We aim to foster a transparent and informed discussion about the use of online video data in computational research.